A random sample of 8 cigarettes of a certain brand
1
Expert's answer
2020-06-30T07:54:25-0400

let the average nicotine content be 4.2 milligrams and a standard deviation - sd=1.4 milligrams.

H0 : μ=3.5 against H1 : μ>3.5

T=(X-μ)/(sd/"\sqrt n") =1.414

table value for 1% significance and df=6 equals to 3.143>1.414, so we accept H0